OpenStudy (kirbykirby):

(Linear Algebra): \[Let: \mathbb{S}=Span \left\{ \left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right],\left[\begin{matrix}0 & 1 \\ 1 & 0\end{matrix}\right],\left[\begin{matrix}1 & -1 \\ 0 & 1\end{matrix}\right] \right\}\] Consider \[A=\left[\begin{matrix}0 & 2 \\ -1 & 2\end{matrix}\right]\]. Find the matrix B in S such that ||A - B|| is minimized.

OpenStudy (kirbykirby):

We saw something "similar" in class but it involved minimizing systems ||Ax-b|| (where x, b are vectors). I'm not sure how it works with just matrices...

OpenStudy (cruffo):

I recall that ||A(x,y)||^2 = (0x+2y)^2 + (-1x + 2y)^2

OpenStudy (cruffo):

so probably what you were doing in class...

OpenStudy (cruffo):

maybe let B = (a,b) (c,d) so C = A-B = (0-a, 2-b) (-1-c, 2-d) then minimize the functional f(x,y) = ||C(x,y)|| problem is that would include 4 unknows.... not nice...

OpenStudy (cruffo):

actually since B is in S, then B is a linear combination of the 3 matrices, so only 3 unknowns: \[ B = a\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right] +b\left[\begin{matrix}0 & 1 \\ 1 & 0\end{matrix}\right] +c\left[\begin{matrix}1 & -1 \\ 0 & 1\end{matrix}\right]\] \[ \quad = \left[\begin{matrix}a & 0 \\ 0 & a\end{matrix}\right] +\left[\begin{matrix}0 & b \\ b & 0\end{matrix}\right] +\left[\begin{matrix}c & -c \\ 0 & c\end{matrix}\right]\] \[ \quad = \left[\begin{matrix}a+c & b-c \\ b & a+c\end{matrix}\right]\]

OpenStudy (cruffo):

Define \[C := A - B \] \[\quad =\left[\begin{matrix}0 & 2 \\ -1 & 2\end{matrix}\right] - \left[\begin{matrix}a+c & b-c \\ b & a+c\end{matrix}\right]\] \[ \quad = \left[\begin{matrix}-a-c & 2-b+c \\-1- b & 2-a-c\end{matrix}\right]\]